OpenStudy (kkutie7):

Please make sure I'm doing this right. \[\int\frac{3y^{3}+5y-1}{y^{3}+y}dy\]

OpenStudy (kkutie7):

\[\rightarrow \int\frac{3y^{3}+5y-1}{y(y^{2}-1)}dy\]

OpenStudy (kkutie7):

\[\rightarrow\int\frac{3y^{3}+5y-1}{y(y-1)(y+1)}dy\] \[\rightarrow \frac{3y^{3}+5y-1}{y(y-1)(y+1)}=\frac{A}{y}+\frac{B}{y-1}+\frac{C}{y+1}\]

OpenStudy (kkutie7):

\[\rightarrow 3y^{3}+5y-1=A(y-1)(y+1)+B(y)(y+1)+C(y)(y-1)\]

OpenStudy (anonymous):

Using partial fraction decomposition for this program will be messy. I haven't worked out all the details but here's what I have so far... I'm pretty sure all of my steps are legal... just breaking up the numerator more and more until we get rid of y^3 altogether.

OpenStudy (kkutie7):

\[\rightarrow x=0\rightarrow -1=-A, A=1\] \[\rightarrow x=1,3+5-1=2B, B=\frac{7}{2}\] \[\rightarrow x=-1,2C, C=\frac{-1}{2} \] I don't see how things are messy, unless I'm doing it wrong

OpenStudy (kkutie7):

\[\rightarrow \int\frac{1}{y}+\frac{7}{2}\int\frac{1}{y-1}+\frac{-1}{2}\int\frac{1}{y+1}\] \

OpenStudy (kkutie7):

\[\rightarrow ln|y|+\frac{7}{2}ln|y-1|-\frac{1}{2}ln|y+1|+C\] is this right?

jimthompson5910 (jim_thompson5910):

A = 1 ... agree B = 7/2 ... agree C = -1/2 ... disagree

OpenStudy (kkutie7):

C=-9/2?

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (kkutie7):

damn ok hold on

ganeshie8 (ganeshie8):

It is all wrong. You must do long division first and make the degree of numerator less than the degree of denominator..

jimthompson5910 (jim_thompson5910):

@ganeshie8 no it's not all wrong. The only mistake I see is that C is the only thing that's off. Everything else looks fine

OpenStudy (kkutie7):

so: